PLS EXPLAIN GREATEST INT FUCTION AND HOW TO DRAW GRAPH OF IT IN DETAIL.


ALSO, PLS DRAW THE GRAPH OF f={(x,x2/1+x2):x belongs toR} FROM R TO R.

Grade:11

1 Answers

Jitender Singh IIT Delhi
9 years ago
Ans:
f(x) = [x]
[ ]represent greatest integer function.
[ 2.3 ] = 2
[ -3.4 ] = -4
[x] = 0, 0\leq x< 1
= 1, 1\leq x< 2
= 2, 2\leq x< 3
Just like that for all the intervals b/w two consecutive integers. You can draw the graph easily now.
Thanks & Regards
Jitender Singh
IIT Delhi
askIITians Faculty